# arXiv:2603.01030v1[math.NA] 01 Mar 2026

### 01 Mar 2026 v1

@@ Abstract
This paper studies pressure-robustness for the axisymmetric Stokes problem. The transformation to cylindrical coordinates requires that the radially weighted velocity is divergence-free in the classical sense. Consequently, traditional divergence-free finite element methods from the Cartesian setting – even if inf–sup stable – are in general not divergence-free in the axisymmetric formulation.
We therefore explore the approach that restores pressure-robustness via reconstruction operators for a low-order Bernardi–Raugel discretization. We show that an application of standard interpolation operators from the Cartesian setting to radially weighted test functions works in principle, but it lacks properties needed to derive optimal consistency error estimates.
@@ 1 Introduction
Another difficulty stems from the $r^{-1}$ weights in the bilinear form $a(\cdot,\cdot)$: they require that the radial component $u_{r}$ vanish at the rotation axis, which is natural for radially symmetric flows but complicates discretizations with $H(\mathrm{div})$-conforming ansatz spaces. To the best of the authors’ knowledge, only a few exactly divergence-free discretizations on structured meshes are available [BFO96, LZZZ].
@@
All considered reconstruction operators $\Pi$ map discretely divergence-free test functions to exactly divergence-free $H(\mathrm{div})$-conforming ones by utilizing the commutative property $\mathrm{div}_{(r,z)}\Pi(r\bm{v}_{h})=\pi_{0}\mathrm{div}_{(r,z)}(r\bm{v}_{h})$. Two approaches are discussed. The first one employs classical $H(\mathrm{div})$-conforming standard interpolations, that are only guaranteed to work under higher regularity assumptions on the problem data. To fix this flaw, a second approach employs a modified $H(\mathrm{div})$-conforming finite element space that guarantees that also the tangential component of the reconstructed functions $r\bm{v}_{h}$ vanish at the rotation axis. This property allows $L^{2}_{-1}$ estimates of the reconstructions and is the key for an improved and optimal consistency error estimate. The theoretical results and the properties of the modified reconstruction operators are confirmed in several numerical experiments. Moreover, the reconstruction operator applied to the discrete velocity solution is divergence-free (in the sense above) and can be used in coupled axisymmetric transport simulations to guarantee maximum principles and mass conservation [FMR26].
Furthermore, the reconstruction space motivates $H(\mathrm{div})$-conforming (H)-DG in the spirit of [MR2304270, LS16, MR4122492, MR3833698]. A thorough investigation of this possibility is the topic of a forthcoming paper.
@@ 2.2 Strong formulation and differential operators
$\Delta_{(r,z)}v:=\partial^{2}_{r}v+\partial^{2}_{z}v,\quad\nabla_{(r,z)}\bm{v}:=(\partial_{r}v_{r},\partial_{z}v_{z}),\quad\text{and}\quad\mathrm{div}_{(r,z)}\bm{v}:=\partial_{r}v_{r}+\partial_{z}v_{z},$
@@ 2.4 Helmholtz projection
$(\nabla\alpha_{\bm{f}},\bm{v}_{0})_{L^{2}_{1}}=(\nabla\alpha_{\bm{f}},r\bm{v}_{0})_{L^{2}}=(\alpha_{\bm{f}},\mathrm{div}_{(r,z)}(r\bm{v}_{0}))_{L^{2}}=0.$
@@ 3.1 A modified finite element method
$\bm{V}_{h}:=\bm{P}_{1}(\mathcal{T})\cup\{\lambda_{i}\lambda_{j}\bm{n}_{E}:E=\mathrm{conv\{P_{i},P_{j}\}}\in\mathcal{E}\},\qquad Q_{h}:=P_{0}(\mathcal{T})\cap L^{2}_{1,0}(\Omega).$
Here, $\lambda_{j}$ is the barycentric coordinate for the vertex $P_{j}$ and $\bm{n}_{E}$ is a fixed unit normal vector of the edge $E\in\mathcal{E}$. The space of discretely divergence-free functions is then given by
@@ 3.2 A priori error estimates
Apart from the Helmholtz decomposition of the right-hand side $\bm{f}$ from (8), the analysis utilizes the discrete Stokes projector $\Pi_{S}$ and the continuous Stokes lifting $\mathcal{L}_{S}$ defined via
@@
$\|\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h}\|_{L^{2}_{1}}\lesssim h^{s}\|\operatorname{div}(r\bm{w}_{h})\|_{L^{2}_{-1}}\lesssim h^{s}\|\bm{w}_{h}\|_{\bm{V}}\quad\text{for all }\bm{w}_{h}\in\bm{V}_{h}.$
@@
$\|\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h}\|_{L^{2}_{1}}^{2}$ $=a(\bm{\psi},\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h})+b(\lambda,\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h})$
$=b(\lambda,\Pi_{S}\bm{w}_{h})$
$=b(\lambda-\pi_{Q_{h}}\lambda,\Pi_{S}\bm{w}_{h})$
@@
This and a similar estimate as in (5) concludes the proof.
@@
$\|\bm{u}-\bm{u}_{h}\|_{\bm{V}}\lesssim\inf_{\bm{v}_{h}\in\bm{V}_{h}^{0}}\|\bm{u}-\bm{v}_{h}\|_{\bm{V}}+h^{s}\|\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}+\mathcal{R}_{\Pi}.$
@@
By definition of the discrete Stokes projector, the error can be split into a best-approximation error and a consistency error via the Pythagorean identity
$\|\bm{u}-\bm{u}_{h}\|^{2}_{\bm{V}}$ $=\inf_{\bm{v}_{h}\in\bm{V}_{h}}\|\bm{u}-\bm{v}_{h}\|^{2}_{\bm{V}}+\|\Pi_{S}\bm{u}-\bm{u}_{h}\|^{2}_{\bm{V}}.$
Since $\bm{w}_{h}:=\Pi_{S}\bm{u}-\bm{u}_{h}\in\bm{V}_{h,0}$ is discretely divergence-free, one obtains
$a(\bm{u}_{h},\bm{w}_{h})$ $=\frac{1}{\nu}(\bm{f},\Pi(r\bm{w}_{h}))_{L^{2}}=\frac{1}{\nu}(\mathbb{P}\bm{f},\Pi(r\bm{w}_{h}))_{L^{2}}$
@@
(17) $a(\bm{u}_{h},\bm{w}_{h})=(\mathbb{P}(-\Delta_{\text{axi}}\bm{u}),\Pi(r\bm{w}_{h})-r\bm{w}_{h})_{L^{2}}+(\mathbb{P}(-\nu\Delta_{\text{axi}}\bm{u}),\bm{w}_{h})_{L^{2}_{1}}\\ +\nu^{-1}(\nabla\alpha_{\bm{f}},\Pi(r\bm{w}_{h}))_{L^{2}}.$
@@
$(\mathbb{P}(-\nu\Delta_{\text{axi}}\bm{u}),\bm{w}_{h})_{L^{2}_{1}}$ $=(\mathbb{P}(-\nu\Delta_{\text{axi}}\bm{u}),\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h})_{L^{2}_{1}}+(\mathbb{P}(-\nu\Delta_{\text{axi}}\bm{u}),\mathcal{L}_{S}\bm{w}_{h})_{L^{2}_{1}}$
(18) $=(\mathbb{P}(-\nu\Delta_{\text{axi}}\bm{u}),\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h})_{L^{2}_{1}}+a(\Pi_{S}\bm{u},\bm{w}_{h}),$
@@
$\leq\left(\|\mathbb{P}(\Delta_{\text{axi}}\bm{u})\circ(\mathbb{I}-\Pi)\|_{\bm{V}_{h,0}^{\star}}+h^{s}\|\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}+\nu^{-1}\|\nabla\alpha_{\bm{f}}\circ\Pi\|_{\bm{V}_{h,0}^{\star}}\right)\|\bm{w}_{h}\|_{V},$
@@ 4 Classical $H(\mathrm{div}$)-conforming reconstruction operators
If the pressure space $Q_{h}$ consists of piecewise discontinuous polynomials of order k1k-1, the standard interpolation operators $\Pi_{\mathrm{BDM1}}$ or $\Pi_{\mathrm{RT0}}$ into the Brezzi–Douglas–Marini (BDM) finite element space
$\mathrm{BDM}_{k}(\mathcal{T}):=P_{k}(\mathcal{T})\cap H(\mathrm{div},\Omega),$
@@
$\mathrm{RT}_{k-1}(\mathcal{T}):=\Big\{\bm{v}_{h}\in$ $P_{k}(\mathcal{T})\cap\mathrm{BDM}_{k}(\mathcal{T}):$
$\forall T\in\mathcal{T}\penalty 10000\ \exists a,b,c\in P_{k-1}(T):\bm{v}_{h}(r,z)|_{T}=\begin{pmatrix}a\\ b\end{pmatrix}+c\begin{pmatrix}r\\ z\end{pmatrix}\Big\}$,
@@
Besides the regularity issue above, the consistency error in Theorem 3.3 can be estimated for the standard reconstruction operators under slightly larger regularity assumptions on the data.
@@
$\|\mathbb{P}(\Delta_{\text{axi}}\bm{u})\circ(\mathbb{I}-\Pi)\|_{\bm{V}^{\star}}\lesssim\|h_{\mathcal{T}}\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}}$
@@
$(\mathbb{P}(\Delta_{\text{axi}}\bm{u}),r\bm{w}_{h}-\Pi(r\bm{w}_{h}))_{L^{2}}$ $\leq\|h_{\mathcal{T}}\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}}\|h_{\mathcal{T}}^{-1}(1-\Pi)(r\bm{w}_{h})\|_{L^{2}}$
$\lesssim\|h_{\mathcal{T}}\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}}\|\bm{w}_{h}\|_{\bm{V}}.$
@@ 5.1 Modified $L^{2}_{-1}(\Omega)\cap H(\mathrm{div},\Omega)$-conforming interpolation
collecting all interior edges that have one vertex on the rotation axis. In Figure 3, we show that there are two possible types of triangles appearing on the rotation axis, which require special care in some parts of the analysis for weighted Sobolev norms.
@@
Figure 3: We distinguish between type 2 ($\text{T}2$) with two vertices on the rotation axis $\Gamma_{\!\text{rot}}$ and type 1 ($\text{T}1$) triangles with one vertex on the rotation axis $\Gamma_{\!\text{rot}}$. Both types of triangles have two edges belonging to $\mathcal{E}_{R}$ on which we define a modified basis function $\psi_{E}^{\mathrm{R}}$ in (26).
@@
Consider an edge $E=\mathrm{conv}\{P_{i},P_{j}\}\in\mathcal{E}_{R}$ with vertex $P_{j}=(r_{j},z_{j})\in\Gamma_{\!\text{rot}}$ on the rotation axis (recall that both type 1 and type 2 triangles have two of them). For this edge $E$, we define a modified $\mathrm{RT}_{0}$ function by
@@
where $\lambda_{i}$ and $\lambda_{j}$ are the standard barycentric coordinates associated to the vertices $P_{i}$ and $P_{j}$, respectively. It can be obtained by a linear combination of the standard $\mathrm{RT}_{0}$ and $\mathrm{BDM}_{1}$ basis functions which (up to scaling factors) usually have the form
@@
The function (26) for an edge $E=\mathrm{conv}\{P_{i},P_{j}\}\in\mathcal{E}_{R}$ with $P_{j}\in\Gamma_{\!\text{rot}}$ and $P_{i}\notin\Gamma_{\!\text{rot}}$ has the following properties:
@@
The first three properties follow from straightforward calculations. For property (iv), let $z_{j}$ be the $z$-coordinate of the vertex $P_{j}\in\Gamma_{\!\text{rot}}$. Then the value $\lambda_{i}$ at a point $(r,z)$ can be expressed as $\lambda_{i}=\nabla\lambda_{i}\cdot(r,z-z_{j})$, because $\lambda_{i}$ is zero at the vertex $P_{j}$. This gives
@@
The last integral can be transformed with the transformation $(r,z-z_{j})=\Phi(r,\alpha)=(r,r\sin(\alpha))$ and $\mathrm{det}(D\Phi)=r\cos(\alpha)$, where $\alpha\in[\alpha_{\text{min}},\alpha_{\text{max}}]$ is the angle between the $r$-axis and the line connecting the point $(r,z)\in T$ with the vertex $P_{j}$. Then we obtain
@@
where linear independency follows from property (i) and (ii) of Lemma 5.1. Further, we define the space spans the space

### 13 Mar 2026 v2

# remarkRemark

@@ Abstract
This paper studies pressure-robustness for the axisymmetric Stokes problem. The transformation to cylindrical coordinates requires that the radially weighted velocity describing the mass flux is divergence-free in the classical sense. Consequently, traditional divergence-free finite element methods from the Cartesian setting – even if inf–sup stable – are in general not divergence-free in the axisymmetric formulation.
We therefore explore the approach that restores pressure-robustness via reconstruction operators for a low-order Bernardi–Raugel discretization. We show that an application of standard interpolation operators from the Cartesian setting to radially weighted test functions works in principle, but lacks properties needed to derive optimal consistency error estimates.
@@ 1 Introduction
Another difficulty stems from the $r^{-1}$ weights in the bilinear form $a(\cdot,\cdot)$: they require that the radial component $u_{r}$ vanish at the rotation axis, which is natural for radially symmetric flows but complicates discretizations with $H(\mathrm{div})$-conforming ansatz spaces. To the best of the authors’ knowledge, only a few exactly divergence-free axisymmetric discretizations on structured meshes are available [BFO96, LZZZ].
@@
All considered reconstruction operators $\Pi$ map discretely divergence-free (radially weighted) test functions to exactly divergence-free $H(\mathrm{div})$-conforming ones by utilizing the commutative property $\mathrm{div}_{(r,z)}\Pi(r\bm{v}_{h})=\pi_{0}\mathrm{div}_{(r,z)}(r\bm{v}_{h})$. Two approaches are discussed. The first one employs classical $H(\mathrm{div})$-conforming standard interpolations, that are only guaranteed to work under higher regularity assumptions on the problem data. To fix this flaw, a second approach employs a modified $H(\mathrm{div})$-conforming finite element space that guarantees that also the tangential component of the reconstructed functions $r\bm{v}_{h}$ vanish at the rotation axis. This property allows $L^{2}_{-1}$ estimates of the reconstructions (where 1-1 indicates the scaling of the radial weight in the $L^{2}$ inner product) and is the key for an improved and optimal consistency error estimate. The theoretical results and the properties of the modified reconstruction operators are confirmed in several numerical experiments. Moreover, the reconstruction operator applied to the discrete velocity solution is divergence-free (in the sense above) and can be used in coupled axisymmetric transport simulations to guarantee maximum principles and mass conservation [FMR26].
Furthermore, the reconstruction space motivates $H(\mathrm{div})$-conforming (H)DG in the spirit of [MR2304270, LS16, MR4122492, MR3833698]. A thorough investigation of this possibility is the topic of a forthcoming paper.
@@ 2.2 Strong formulation and differential operators
$\Delta_{(r,z)}v:=\partial^{2}_{r}v+\partial^{2}_{z}v,\quad\nabla_{(r,z)}\bm{v}\cdot\bm{e}_{s}:=\partial_{s}\bm{v},s\in\{r,z\},\text{ and }\mathrm{div}_{(r,z)}\bm{v}:=\partial_{r}v_{r}+\partial_{z}v_{z},$
@@ 2.4 Helmholtz projection
$(\nabla\alpha_{\bm{f}},\bm{v}_{0})_{L^{2}_{1}}=(\nabla\alpha_{\bm{f}},r\bm{v}_{0})_{L^{2}}=-(\alpha_{\bm{f}},\mathrm{div}_{(r,z)}(r\bm{v}_{0}))_{L^{2}}=0.$
@@ 3.1 A modified finite element method
$\bm{V}_{h}:=\bm{P}_{1}(\mathcal{T})\oplus\{\lambda_{i}\lambda_{j}\bm{n}_{E}:E=\mathrm{conv}\{N_{i},N_{j}\}\in\mathcal{E}\},\qquad Q_{h}:=P_{0}(\mathcal{T})\cap L^{2}_{1,0}(\Omega).$
Here, $\lambda_{j}$ is the barycentric coordinate for the vertex $N_{j}$ and $\bm{n}_{E}$ is a fixed unit normal vector of the edge $E\in\mathcal{E}$. The space of discretely divergence-free functions is then given by
@@ 3.2 A priori error estimates
Apart from the Helmholtz decomposition of the right-hand side $\bm{f}$ from (8), the analysis utilizes the discrete Stokes projector $\Pi_{S}:\bm{V}\to\bm{V}_{h,0}$ and the continuous Stokes lifting $\mathcal{L}_{S}:\bm{V}_{h}\to\bm{V}_{0}$ defined via
@@
(14) $\|\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h}\|_{L^{2}_{1}}\lesssim h^{s}\|\operatorname{div}(r\bm{w}_{h})\|_{L^{2}_{-1}}\lesssim h^{s}\|\bm{w}_{h}\|_{\bm{V}}\quad\text{for all }\bm{w}_{h}\in\bm{V}_{h,0}.$
@@
$\|\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h}\|_{L^{2}_{1}}^{2}$ $=a(\bm{\psi},\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h})+b(\lambda,\bm{w}_{h})-b(\lambda,\mathcal{L}_{S}\bm{w}_{h})=0+b(\lambda,\bm{w}_{h})+0$
$=b(\lambda-\pi_{Q_{h}}\lambda,\bm{w}_{h})\qquad{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}(\text{as }b(q_{h},\bm{w}_{h})=0\text{ for all }q_{h}\in Q_{h})}$
@@
Following the estimates in (5) concludes the proof.
@@
$\|\bm{u}-\bm{u}_{h}\|_{\bm{V}}\lesssim\inf_{\bm{v}_{h}\in\bm{V}_{h,0}}\|\bm{u}-\bm{v}_{h}\|_{\bm{V}}+h^{s}\|\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}+\mathcal{R}_{\Pi}.$
@@
We first note that $\bm{u}\in\bm{V}_{0}$ and $\bm{u}_{h}\in\bm{V}_{h,0}$ and by definition of the discrete Stokes projector we have $\bm{w}_{h}:=\Pi_{S}\bm{u}-\bm{u}_{h}\in\bm{V}_{h,0}$ and $a(\Pi_{S}\bm{u}-\bm{u},\Pi_{S}\bm{u}-\bm{u}_{h})=0$, i.e. $\Pi_{S}\bm{u}-\bm{u}\perp_{a}\Pi_{S}\bm{u}-\bm{u}_{h}$. Hence, the error can be split into a best-approximation error and a consistency error via the Pythagorean identity
$\|\bm{u}-\bm{u}_{h}\|^{2}_{\bm{V}}$ $=\|\bm{u}-\Pi_{S}\bm{u}\|^{2}_{\bm{V}}+\|\Pi_{S}\bm{u}-\bm{u}_{h}\|^{2}_{\bm{V}}.$
where $\|\bm{u}-\Pi_{S}\bm{u}\|^{2}_{\bm{V}}=\inf_{\bm{v}_{h}\in\bm{V}_{h,0}}\|\bm{u}-\bm{v}_{h}\|^{2}_{\bm{V}}$. It hence remains only to estimate the consistency error second term $\|\Pi_{S}\bm{u}-\bm{u}_{h}\|^{2}_{\bm{V}}$. As $b(\bm{w}_{h},q_{h})=0$ for all $q_{h}\in Q_{h}$, we further obtain
$a(\bm{u}_{h},\bm{w}_{h})$ $=\nu^{-1}(\bm{f},\Pi(r\bm{w}_{h}))_{L^{2}}=\nu^{-1}(\mathbb{P}\bm{f}+\nabla\alpha_{\bm{f}},\Pi(r\bm{w}_{h}))_{L^{2}}$
@@
(18) $a(\bm{u}_{h},\bm{w}_{h})=(\mathbb{P}(-\Delta_{\text{axi}}\bm{u}),\Pi(r\bm{w}_{h})-r\bm{w}_{h})_{L^{2}}+(\mathbb{P}(-\Delta_{\text{axi}}\bm{u}),\bm{w}_{h})_{L^{2}_{1}}\\ +\nu^{-1}(\nabla\alpha_{\bm{f}},\Pi(r\bm{w}_{h}))_{L^{2}}.$
@@
$(\mathbb{P}(-\Delta_{\text{axi}}\bm{u}),\bm{w}_{h})_{L^{2}_{1}}$ $=(\mathbb{P}(-\Delta_{\text{axi}}\bm{u}),\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h})_{L^{2}_{1}}+(\mathbb{P}(-\Delta_{\text{axi}}\bm{u}),\mathcal{L}_{S}\bm{w}_{h})_{L^{2}_{1}}$
(19) $=(\mathbb{P}(-\Delta_{\text{axi}}\bm{u}),\bm{w}_{h}-\mathcal{L}_{S}\bm{w}_{h})_{L^{2}_{1}}+a(\Pi_{S}\bm{u},\bm{w}_{h}),$
@@
[…]
@@ 4 Classical $H(\mathrm{div}$)-conforming reconstruction operators
If the pressure space $Q_{h}$ consists of piecewise discontinuous polynomials of order 0, the standard interpolation operators $\Pi_{\mathrm{BDM1}}$ or $\Pi_{\mathrm{RT0}}$ into the Brezzi–Douglas–Marini (BDM) finite element space
$\mathrm{BDM}_{1}(\mathcal{T}):=\bm{P}_{1}(\mathcal{T})\cap H(\mathrm{div},\Omega),$
@@
$\mathrm{RT}_{0}(\mathcal{T}):=\Big\{\bm{v}_{h}(r,z)|_{T}=\begin{pmatrix}a\\ b\end{pmatrix}+c\begin{pmatrix}r\\ z\end{pmatrix}:a,b,c\in\mathbb{R}\text{ for all }T\in\mathcal{T}\Big\}\cap H(\mathrm{div},\Omega),$
@@
Besides the regularity issue above, the consistency error in Theorem 3.3 can be estimated for the standard reconstruction operators under slightly stronger regularity assumptions on the data.
@@
$\|\mathbb{P}(\Delta_{\text{axi}}\bm{u})\circ(\mathbb{I}-\Pi)\|_{\bm{V}^{\star}}\lesssim\|h\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}}$
@@
$(\mathbb{P}(\Delta_{\text{axi}}\bm{u}),r\bm{w}_{h}-\Pi(r\bm{w}_{h}))_{L^{2}}$ $\leq\|h\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}}\|h^{-1}(1-\Pi)(r\bm{w}_{h})\|_{L^{2}}$
$\lesssim\|h\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}}\|\bm{w}_{h}\|_{\bm{V}}.$
@@ 5.1 Modified $L^{2}_{-1}(\Omega)\cap H(\mathrm{div},\Omega)$-conforming interpolation
collecting all interior edges that have one vertex on the rotation axis. We distinguish two types of triangles adjacent to the rotation axis (see Figure 3): a type 1 triangle has exactly one vertex on $\Gamma_{\!\text{rot}}$, while a type 2 triangle has exactly two vertices on $\Gamma_{\!\text{rot}}$. Both types have two edges belonging to $\mathcal{E}_{R}$ and require special care in some parts of the analysis for weighted Sobolev norms.
@@
Figure 3: Illustration of type 2 ($\text{T}2$) and type 1 ($\text{T}1$) triangles adjacent to the rotation axis $\Gamma_{\!\text{rot}}$. Both types of triangles have two edges belonging to $\mathcal{E}_{R}$ on which we define a modified basis function $\psi_{E}^{\mathrm{R}}$ in (27).
@@
Consider an edge $E=\mathrm{conv}\{N_{i},N_{j}\}\in\mathcal{E}_{R}$ with vertex $N_{j}=(r_{j},z_{j})\in\Gamma_{\!\text{rot}}$ on the rotation axis (recall that both type 1 and type 2 triangles have two of them). For this edge $E$, we define a modified $\mathrm{RT}_{0}$ function by
@@
where $\lambda_{i}$ and $\lambda_{j}$ are the standard barycentric coordinates associated to the vertices $N_{i}$ and $N_{j}$, respectively. It can be obtained by a linear combination of the standard $\mathrm{RT}_{0}$ and $\mathrm{BDM}_{1}$ basis functions which (up to scaling factors) usually have the form
@@
The function (27) for an edge $E=\mathrm{conv}\{N_{i},N_{j}\}\in\mathcal{E}_{R}$ with $N_{j}\in\Gamma_{\!\text{rot}}$ and $N_{i}\notin\Gamma_{\!\text{rot}}$ has the following properties:
@@
The first three properties follow from straightforward calculations. For property (iv), let $z_{j}$ be the $z$-coordinate of the vertex $N_{j}\in\Gamma_{\!\text{rot}}$. Then the value $\lambda_{i}$ at a point $(r,z)$ can be expressed as $\lambda_{i}=\nabla\lambda_{i}\cdot(r,z-z_{j})$, because $\lambda_{i}$ is zero at the vertex $N_{j}$. This gives
@@
The last integral can be transformed with the transformation $(r,z-z_{j})=\Phi(r,\alpha)=(r,r\sin(\alpha))$ and $\mathrm{det}(D\Phi)=r\cos(\alpha)$, where $\alpha\in[\alpha_{\text{min}},\alpha_{\text{max}}]$ is the angle between the $r$-axis and the line connecting the point $(r,z)\in T$ with the vertex $N_{j}$. Then we obtain
@@
where linear independence follows from property (i) and (ii) of Lemma 5.1. Further, we define the space